OpenStudy (anonymous):

jacob scored 5 points more than 4/5 of Sarai's test score. which equation represents the relationship between Jacob's test score, J, and Sarai's test score, S?

OpenStudy (beccaboo333):

Are there options?

OpenStudy (anonymous):

yes there is F. 4/5 J+ 5 =S G. 4/5 s + 5 =J H. 4/5 J - 5 = S I. 4/5 S - 5 = J

OpenStudy (beccaboo333):

4/5 s + 5 =J I don't really know how to explain it sorreh

OpenStudy (anonymous):

Its fine.